Circular No. 6863

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1998 G1 (LINEAR) Of the many fast-moving asteroidal objects recently reported by the

Lincoln Laboratory Near-Earth Asteroid Research project, one has a retrograde, nearly-parabolic orbit. In response to our request, W. B. Offutt coadded 12 CCD frames (appropriately shifted) of the object and notes that “there appears to be a clear and unmistakable tail in p.a. 95^{o}, traceable out to 19′′ from the center of the nucleus; the tail is fairly well defined and straight; the nucleus remains hard and well defined, and I detect no coma”.

1998 UT α_{2000} δ_{2000} m_{2} Observer
Apr. 2.13056 11^{h}08^{m}19.^{s}60 +29^{o}36′13′′.2 18.9 Blythe
3.13816 11 05 47.52 +29 29 31.1 18.3  _{”}
3.54554 11 04 46.20 +29 26 45.5 17.4 Garradd
4.51542 11 02 21.69 +29 19 48.9 18.0 Zoltowski
4.54472 11 02 17.38 +29 19 34.0 17.9 Sugie
4.57462 11 02 12.84 +29 19 20.4 17.9 Ikari
5.13988 11 00 49.64 +29 15 07.8 17.6 Offutt
5.51900 10 59 54.01 +29 12 18.2  Garradd
5.80000 10 59 12.90 +29 10 02.8 18.4 Šarounová
6.08351 10 58 31.60 +29 07 50.6 17.7 DeGraff

M. Blythe, F. Shelly and M. Bezpalko (Lincoln Laboratory ETS). 1-m

f/2.15 reflector. Measurers J. Stuart and H. Viggh. G. J. Garradd (Loomberah). 0.25-m f/4.1 reflector. F. Zoltowski (Woomera). 0.30-m f/3.3 Schmidt-Cassegrain. A. Sugie (Dynic Astronomical Observatory). 0.60-m f/5.0 reflector. Y. Ikari (Moriyama). 0.25-m f/6.3 Schmidt-Cassegrain. W. B. Offutt (Cloudcroft). 0.60-m f/7 Ritchey-Chrétien. L. Šarounová (Ondřejov). 0.65-m f/3.6 reflector. D. R. DeGraff and A. C. Schwortz (Alfred University). 0.81-m f/4 reflector.

T = 1998 Oct. 15.550 TT  ω = 218.^{o}101
Ω = 339.322 }2000.0
q = 2.55877 AU  i = 114.115
1998TT α_{2000} δ_{2000} Δ r ϵ β m_{2}
Apr. 7 10^{h}56.^{m}32 +29^{o}00.^{′}5 2.489 3.256 133.^{o}1 13.^{o}0 18.3
12 10 44.76 +28 15.7 2.517 3.225 127.3 14.3 18.4
17 10 34.11 +27 24.0 2.555 3.194 121.3 15.6 18.4
22 10 24.47 +26 26.8 2.600 3.164 115.4 16.7 18.5
27 10 15.87 +25 25.5 2.652 3.134 109.5 17.6 18.5
1998 April 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden